# thanks: M.C. and D.P. contributed equally to this work

### v1

@@ Abstract
Primordial black holes (PBHs) are usually assumed to be described by the Schwarzschild or Kerr metrics, which however feature unwelcome singularities. We study the possibility that PBHs are non-singular objects, considering four phenomenological, regular tr-symmetric space-times, featuring either de Sitter or Minkowski cores. We characterize the evaporation of these PBHs and constrain their abundance from $\gamma$-ray observations. For three of the metrics, including the well-known Bardeen and Hayward ones, we show that constraints on $f_{\text{pbh}}$, the fraction of dark matter (DM) in the form of PBHs, weaken with respect to the Schwarzschild limits, because of modifications to the PBH temperature and greybody factors. This moves the lower edge of the asteroid mass window down by up to an order of magnitude, leading to a much larger region of parameter space where PBHs can make up all the DM. A companion paper is instead devoted to non-tr-symmetric metrics, including loop quantum gravity-inspired ones. Our work provides a proof-of-principle for the interface between the DM and singularity problems being a promising arena with a rich phenomenology.
@@ I Introduction
The Standard Model of particle physics (SM) and General Relativity (GR) have proven to be extremely successful at describing a huge range of terrestrial, astrophysical, and cosmological observations Gaillard et al. (1999); Will (2014). However, their successes are limited by a number of shortcomings, potentially (especially in the case of SM) pointing towards the need for new physics which may better describe the matter and gravity sectors. On the more observational/phenomenological side, the SM lacks a candidate for the dark matter (DM) which accounts for $\simeq 25\%$ of the energy budget of the Universe Arbey and Mahmoudi (2021); Cirelli et al. (2024). On the more theoretical side, continuous gravitational collapse in GR leads to the pathological appearance of curvature singularities Penrose (1965); Hawking and Penrose (1970). The nature of DM and the singularity problem are arguably two among the most important open questions in theoretical physics.
@@
The solution to the former problem could reside in the physics of some of the most peculiar objects in the Universe: black holes (BHs). It has long been realized that primordial BHs (PBHs), hypothetical relics from the primordial Universe formed from the collapse of large density perturbations upon horizon re-entry, are indeed excellent DM candidates Chapline (1975); Meszaros (1975); Ivanov et al. (1994); Bird et al. (2016); Clesse and García-Bellido (2017); Poulin et al. (2017); Raccanelli et al. (2018); Bernal et al. (2017); Clesse and García-Bellido (2018); Kohri and Terada (2018); Murgia et al. (2019); Carr et al. (2021a); Serpico et al. (2020); De Luca et al. (2020a, b, c); Carr et al. (2021b); Bhagwat et al. (2021); De Luca et al. (2021a); Wong et al. (2021); Carr et al. (2021c); Domènech et al. (2021); De Luca et al. (2021b); Arbey et al. (2021a); Franciolini et al. (2022a); De Luca et al. (2021c); Cheek et al. (2022a, b); Dvali et al. (2021); De Luca et al. (2022); Bhaumik et al. (2022); Cai et al. (2022); Oguri et al. (2023); Franciolini et al. (2022b); Mazde and Visinelli (2023); Cai et al. (2023); Fu and Wang (2023); Choudhury et al. (2024a, 2023); Musco et al. (2024); Yuan et al. (2024); Ghoshal et al. (2023); Choudhury et al. (2024b); Ghoshal and Strumia (2024); Thoss et al. (2024); Papanikolaou et al. (2024a, b); Yin and Visinelli (2024); Boccia et al. (2024); Yang et al. (2024) (see e.g. Refs. Carr et al. (2016a); Green and Kavanagh (2021); Carr and Kuhnel (2020, 2022); Bird et al. (2023); Carr et al. (2024); Choudhury and Sami (2024) for reviews): in fact, PBHs are the only viable DM candidate which does not invoke new particles surviving to the present day. Once believed to merely be objects of mathematical speculation, observational effects associated to BHs are now routinely detected Bambi (2018), turning these objects into extraordinary probes of fundamental physics Creminelli and Vernizzi (2017); Sakstein and Jain (2017); Ezquiaga and Zumalacárregui (2017); Boran et al. (2018); Baker et al. (2017); Amendola et al. (2018); Visinelli et al. (2018); Crisostomi and Koyama (2018); Dima and Vernizzi (2018); Cai et al. (2018); Casalino et al. (2018); Barack et al. (2019); Abbott et al. (2019); Casalino et al. (2019); Held et al. (2019); Bambi et al. (2019); Vagnozzi and Visinelli (2019); Zhu et al. (2019); Cunha et al. (2019); Banerjee et al. (2020a, b); Allahyari et al. (2020); Vagnozzi et al. (2020); Khodadi et al. (2020); Kumar et al. (2020a); Khodadi and Saridakis (2021); Pantig et al. (2022); Khodadi et al. (2021); Roy et al. (2022); Uniyal et al. (2023); Pantig and Övgün (2023); Ghosh and Afrin (2023); Khodadi and Lambiase (2022); Kumar Walia et al. (2022); Shaikh (2023); Odintsov and Oikonomou (2022); Oikonomou et al. (2022); Pantig (2024); González et al. (2023); Sahoo et al. (2024); Nozari and Saghafi (2023); Uniyal et al. (2024); Filho et al. (2024); Raza et al. (2024); Hoshimov et al. (2024); Chakhchi et al. (2024); Liu et al. (2024a, b); Khodadi et al. (2024); Nojiri and Odintsov (2024). As a result, the possibility of PBHs accounting for the entire DM budget is severely constrained by a wide range of considerations and constraints. The only (not entirely undebated) remaining open window of parameter space where PBHs could make up all the DM is the so-called “asteroid mass window”, roughly for PBH masses $10^{14}\,{\text{kg}}\lesssim M_{\text{pbh}}\lesssim 10^{20}\,{\text{kg}}$ Katz et al. (2018); Smyth et al. (2020); Coogan et al. (2021); Ray et al. (2021); Auffinger (2022); Ghosh and Mishra (2024); Miller et al. (2022); Branco et al. (2023); Bertrand et al. (2023); Tran et al. (2023); Gorton and Green (2024); Dent et al. (2024); Tamta et al. (2024); Tinyakov (2024); Loeb (2024): lighter PBHs would have evaporated fast enough to either have disappeared by now or overproduced $\gamma$-rays in the MeV range, whereas heavier PBHs would have been detected through the microlensing of background stars.
@@
Almost all works on PBHs assume that these are Schwarzschild or Kerr BHs Carr et al. (2016a); Green and Kavanagh (2021); Carr and Kuhnel (2020, 2022); Bird et al. (2023); Carr et al. (2024); Choudhury and Sami (2024). All constraints and considerations on DM potentially being in the form of PBHs are therefore subject to this underlying assumption. The existence of the asteroid mass window, and the extension thereof, is of course no exception. The assumption in question is not at all unreasonable at least from the phenomenological point of view, given that there are at present no signs of tension between astrophysical observations and the Kerr-Newman family of metrics, and more generally the no-hair theorem. Nevertheless, from the theoretical point of view such an assumption might stir some unease, given the appearance of singularities in the Schwarzschild and Kerr metrics. The above considerations naturally lead to the following question: “what if PBHs are non-singular”? It is our goal in the present work to systematically address this question, which naturally merges the DM and singularity problems.
@@
We entertain the possibility that PBHs are “regular”, i.e. free of curvature singularities Ansoldi (2008); Nicolini (2009); Sebastiani and Zerbini (2022); Torres (2022); Lan et al. (2023), and therefore that DM may be in the form of primordial regular BHs (PRBHs). For concreteness, we consider so-called tr (time-radius)-symmetric metrics, for which the product of the coefficients for the $dt^{2}$ and $dr^{2}$ components of the line element in four-dimensional Boyer–Lindquist coordinates is equal to $-1$, and the function which multiplies the angular part of the line element is $r^{2}$, i.e. $r$ is the areal radius. More specifically, we focus on the following four regular, static spherically symmetric space-times, all of which are characterized by an additional regularizing parameter $\ell$ and recover the Schwarzschild space-time in the $\ell\to 0$ limit: Bardeen BHs Bardeen (1968), Hayward BHs Hayward (2006), Ghosh-Culetu-Simpson-Visser BHs Ghosh (2015); Culetu (2015); Simpson and Visser (2019), and a Hayward-like BH which, to the best of our knowledge, is introduced for the first time in this work. These four space-times present a rich phenomenology, including both de Sitter and Minkowski cores, and temperatures which can both decrease and increase with increasing regularizing parameter $\ell$. We focus our attention on observational constraints from PRBH evaporation, which set the lower limit of the asteroid mass window, discussing in detail how the evaporation process is modified with respect to that for Schwarzschild PBHs. We show that, as a result, the phenomenology of PRBHs can be quite different from that of Schwarzschild PBHs, with a larger range of masses where PRBHs could make up the entire DM component, opening up the asteroid mass window by up to an extra decade in mass. Keeping in mind that the metrics in question are phenomenological in nature, our results demonstrate that a common solution to the DM and singularity problems in the form of primordial regular BHs is one which is worth taking seriously, and warrants further investigation, and more generally the interface of these two problems provides a promising arena. We stress that our work should not be intended as a comprehensive analysis of primordial regular BHs, but rather as a pilot study, pointing towards a direction which has thus far received very little attention and indicating promising directions for further work.
@@
The rest of this paper is then organized as follows. In Sec. II we briefly introduce the regular space-times studied in the rest of the work. Various aspects of our methodology are discussed in Sec. III, with Sec. III.1 devoted to the calculation of the so-called greybody factors, Sec. III.2 to the computation of photon spectra resulting from Hawking evaporation, and Sec. III.3 to the comparison against observations. The resulting limits on the fraction of DM which may be in the form of primordial regular BHs are then critically discussed in Sec. IV. Finally, in Sec. V we draw concluding remarks. A number of more technical aspects concerning the greybody factors computation are discussed in Appendix A. Unless otherwise specified, we adopt units where $G=c=1$. In closing, we note that a related study is being presented in a companion paper Calzà et al. (tion): this focuses on non-tr-symmetric metrics, which also include loop quantum gravity-inspired metrics, but at the same time complicate the study of the evaporation process. We recommend that the interested reader go through the present work prior to consulting our companion paper Calzà et al. (tion).
@@ II Regular black holes
It is well known, thanks to the Penrose-Hawking singularity theorems, that continuous gravitational collapse in GR sourced by matter contents satisfying reasonable energy conditions leads to the appearance of singularities Penrose (1965); Hawking and Penrose (1970). These are regions of space-time where curvature invariants, i.e. sets of independent scalars constructed from the Riemann tensor and the metric, diverge (with the archetypal example being the central singularity in the Kerr-Newman family of metrics). These singularities are arguably unsatisfactory as they lead to a potential breakdown in predictivity. For this reason, they are oftentimes regarded as a manifestation of our lack of knowledge of (new) physics in the high-energy/high-curvature regime. A widespread belief is that quantum gravity effects on these scales would ultimately cure the singularity problem (and potentially lead to observable effects), although this is more of a hope supported only by a few first-principles studies Dymnikova (1992); Dymnikova and Galaktionov (2005); Ashtekar and Bojowald (2005); Bebronne and Tinyakov (2009); Modesto et al. (2011); Spallucci and Ansoldi (2011); Perez (2015); Colléaux et al. (2018); Nicolini et al. (2019); Bosma et al. (2019); Jusufi (2023a); Olmo and Rubiera-Garcia (2022); Jusufi (2023b); Ashtekar et al. (2023); Nicolini (2023).
@@
Even in the absence of a widely agreed upon theory of quantum gravity, one can still hope to make progress in understanding and taming singularities, while also potentially gaining intuition about the possible features of such a theory, through a more phenomenological approach. Under the assumption that a metric description holds valid, one can introduce metrics which are free of singularities in the entire space-time, and describe so-called regular BHs (RBHs) Bambi (2023). It is often (albeit not necessarily always) the case that RBH metrics are controlled by an extra parameter, which in what follows we shall refer to as regularizing parameter (and denote by $\ell$), typically recovering the Schwarzschild metric (for non-rotating RBHs) in the limit $\ell\to 0$. Several RBH metrics have been studied over the past decades, see e.g. Refs. Borde (1997); Ayon-Beato and Garcia (1998, 1999); Bronnikov and Fabris (2006); Berej et al. (2006); Bronnikov et al. (2012); Rinaldi (2012); Stuchlík and Schee (2014); Schee and Stuchlik (2015); Johannsen (2013); Myrzakulov et al. (2016); Fan and Wang (2016); Sebastiani et al. (2017); Toshmatov et al. (2017); Chinaglia and Zerbini (2017); Frolov (2018); Bertipagani et al. (2021); Nashed and Saridakis (2022); Simpson and Visser (2022); Franzin et al. (2022); Chataignier et al. (2023); Khodadi and Pourkhodabakhshi (2022); Sajadi et al. (2024); Javed et al. (2024); Ditta et al. (2024); Al-Badawi et al. (2024); Övgün et al. (2024); Corona et al. (2024); Bueno et al. (2024); Pedrotti and Vagnozzi (2024); Konoplya and Zhidenko (2024a); Bronnikov (2024); Bolokhov et al. (2024); Agrawal et al. (2024); Khodadi and Firouzjaee (2024) for an inevitably incomplete selection of examples, as well as Refs. Ansoldi (2008); Sebastiani and Zerbini (2022); Lan et al. (2023) for recent reviews on the subject.  [^1] Another interesting possibility are gravastars, which are not RBHs in a strict sense Mazur and Mottola (2023, 2004); Mottola (2023). While most of these metrics have been introduced on purely phenomenological grounds it is known that possible sources for several RBH metrics lie in theories of non-linear electrodynamics Bronnikov (2018); Ghosh and Walia (2021); Bronnikov and Walia (2022); Bronnikov (2022); Bokulić et al. (2023).
@@
As alluded to earlier, our interest in this work is to explore the possibility that primordial RBHs may play the role of DM. As a proof-of-principle in this sense we will establish constraints on $f_{\text{pbh}}$, the fraction of DM in the form of PRBHs, focusing on the asteroid mass window, whose extent we will show can be either opened or tightened depending on the features of the RBH metric. To the best of our knowledge, we are aware of only four works in this direction Pacheco (2018); Arbey and Auffinger (2021); Arbey et al. (2022); Banerjee et al. (2024). Ref. Pacheco (2018) studied the thermodynamics of primordial regular BHs, focusing however on the case where they do not evaporate, and therefore did not study constraints on $f_{\text{pbh}}$. Ref. Arbey and Auffinger (2021) studied the evaporation of a loop quantum gravity-inspired BH, and weakened constraints on $f_{\text{pbh}}$ were reported in a later proceeding (which however does not appear to be widely known). Finally, Ref. Banerjee et al. (2024) studied signatures of primordial BHs with magnetic charge, which could be (as is often but necessarily the case) regular. The aim of this pilot study and our companion paper Calzà et al. (tion) is instead to provide a more comprehensive investigation of primordial regular BHs, considering a more diverse set of metrics and investigating constraints on $f_{\text{pbh}}$ in detail.
@@
In addition, as stated earlier, we require our space-times to be tr-symmetric (the non-tr-symmetric case is covered in a companion paper Calzà et al. (tion)), which imposes the following additional conditions:
@@
$ds^{2}=-f(r)dt^{2}+f(r)^{-1}dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}(\theta)d\phi^{2}\right)\,.$ (4)
@@
with the first equality following from the choice of focusing on tr-symmetric space-times. In the case of Schwarzschild BHs, where the metric function is $f(r)=1-2M/r$, one recovers the well-known expressions $r_{H}=2M$ and $T_{\text{Sch}}=1/8\pi M$. However, in more general space-times the horizon radius in Eq. (7) is not guaranteed to have a closed form expression, and the same therefore holds for the temperature in Eq. (5). In Fig. 1 we show the evolution of the temperatures (normalized to the temperature of Schwarzschild BHs, $T_{\text{Sch}}=1/8\pi M$) of the four regular space-times we will introduce shortly, as a function of the regularizing parameter $\ell$ (normalized by the event horizon radius $r_{H}$). As the Figure clearly shows, for three of these space-times (Bardeen, Hayward, and Ghosh-Culetu-Simpson-Visser) the temperature is a monotonically decreasing function of $\ell$, whereas for the fourth (Hayward-like), after a small initial dip, the temperature grows monotonically with growing $\ell$. As we will see, these different behaviour turn out to have important phenomenological consequences.

### v2

@@ Abstract
Primordial black holes (PBHs) are usually assumed to be described by the Schwarzschild or Kerr metrics, which however feature unwelcome singularities. We study the possibility that PBHs are non-singular objects, considering four phenomenological, regular tr-symmetric space-times, featuring either de Sitter or Minkowski cores. We characterize the evaporation of these PBHs and constrain their abundance from $\gamma$-ray observations. For three of the metrics, including the well-known Bardeen and Hayward ones, we show that constraints on $f_{\text{pbh}}$, the fraction of dark matter (DM) in the form of PBHs, weaken with respect to the Schwarzschild limits, because of modifications to the PBH temperature and greybody factors. This moves the lower edge of the asteroid mass window down by up to an order of magnitude or more, leading to a much larger region of parameter space where PBHs can make up all the DM. A companion paper is devoted to non-tr-symmetric metrics, including loop quantum gravity-inspired ones. Our work provides a proof-of-principle for the interface between the DM and singularity problems being a promising arena with a rich phenomenology.
@@ I Introduction
The Standard Model of particle physics (SM) and General Relativity (GR) have proven to be extremely successful at describing a huge range of terrestrial, astrophysical, and cosmological observations. However, their successes are limited by a number of shortcomings, potentially (especially in the case of SM) pointing towards the need for new physics which may better describe the matter and gravity sectors. On the more observational/phenomenological side, the SM lacks a candidate for the dark matter (DM) which accounts for $\simeq 25\%$ of the energy budget of the Universe Arbey and Mahmoudi (2021); Cirelli et al. (2024). On the more theoretical side, continuous gravitational collapse in GR leads to the pathological appearance of curvature singularities Penrose (1965); Hawking and Penrose (1970). The nature of DM and the singularity problem are arguably two among the most important open questions in theoretical physics.
@@
The solution to the DM problem could reside in the physics of some of the most peculiar objects in the Universe: black holes (BHs). It has long been realized that primordial BHs (PBHs), hypothetical relics from the primordial Universe formed from the collapse of large density perturbations upon horizon re-entry, are indeed excellent DM candidates Chapline (1975); Meszaros (1975); Khlopov and Polnarev (1980); Khlopov et al. (1985); Ivanov et al. (1994); Choudhury and Mazumdar (2014); Belotsky et al. (2014); Bird et al. (2016); Clesse and García-Bellido (2017); Poulin et al. (2017); Raccanelli et al. (2018); Bernal et al. (2017); Clesse and García-Bellido (2018); Kohri and Terada (2018); Murgia et al. (2019); Carr et al. (2021a); Serpico et al. (2020); De Luca et al. (2020a, b, c); Carr et al. (2021b); Bhagwat et al. (2021); De Luca et al. (2021a); Wong et al. (2021); Carr et al. (2021c); Domènech et al. (2021); De Luca et al. (2021b); Arbey et al. (2021a); Franciolini et al. (2022a); De Luca et al. (2021c); Cheek et al. (2022a, b); Heydari and Karami (2022a); Dvali et al. (2021); Heydari and Karami (2022b); De Luca et al. (2022); Bhaumik et al. (2022); Cai et al. (2022); Oguri et al. (2023); Franciolini et al. (2022b); Mazde and Visinelli (2023); Cai et al. (2023); Fu and Wang (2023); Choudhury et al. (2024a); Papanikolaou (2023); Choudhury et al. (2023a, b); Musco et al. (2024); Yuan et al. (2024); Choudhury et al. (2023c); Ghoshal et al. (2023); Cai et al. (2024a); Choudhury et al. (2024b); Huang et al. (2024a); Choudhury et al. (2023d); Bhattacharya et al. (2024); Heydari and Karami (2024a, b); Choudhury et al. (2024c, 2023e); Ghoshal and Strumia (2024); Huang et al. (2024b); Huang and Piao (2024); Choudhury et al. (2024d); Thoss et al. (2024); Papanikolaou et al. (2024a); Choudhury (2024); Choudhury et al. (2024e); Papanikolaou et al. (2024b); Yin and Visinelli (2024); Choudhury et al. (2024f); Heydari and Karami (2024c); Dvali et al. (2024); Boccia et al. (2024); Huang et al. (2024c); Choudhury et al. (2024g); Yang et al. (2024) (see e.g. Refs. Khlopov (2010); Carr et al. (2016a); Green and Kavanagh (2021); Carr and Kuhnel (2020); Villanueva-Domingo et al. (2021); Carr and Kuhnel (2022); Bird et al. (2023); Carr et al. (2024); Arbey (2024); Choudhury and Sami (2024) for reviews): in fact, PBHs are the only viable DM candidate which does not invoke new particles surviving to the present day. Once believed to merely be objects of mathematical speculation, observational effects associated to BHs are now routinely detected Bambi (2018), turning these objects into extraordinary probes of fundamental physics Creminelli and Vernizzi (2017); Sakstein and Jain (2017); Ezquiaga and Zumalacárregui (2017); Boran et al. (2018); Baker et al. (2017); Amendola et al. (2018); Visinelli et al. (2018); Crisostomi and Koyama (2018); Dima and Vernizzi (2018); Cai et al. (2018); Casalino et al. (2018); Barack et al. (2019); Abbott et al. (2019); Casalino et al. (2019); Held et al. (2019); Bambi et al. (2019); Vagnozzi and Visinelli (2019); Zhu et al. (2019); Cunha et al. (2019); Banerjee et al. (2020a, b); Allahyari et al. (2020); Vagnozzi et al. (2020); Khodadi et al. (2020); Kumar et al. (2020a); Khodadi and Saridakis (2021); Pantig et al. (2022); Khodadi et al. (2021); Roy et al. (2022); Uniyal et al. (2023); Pantig and Övgün (2023); Ghosh and Afrin (2023); Khodadi and Lambiase (2022); Kumar Walia et al. (2022); Shaikh (2023); Odintsov and Oikonomou (2022); Oikonomou et al. (2022); Pantig (2024); González et al. (2023); Sahoo et al. (2024); Nozari and Saghafi (2023); Uniyal et al. (2024); Filho et al. (2024); Raza et al. (2024); Hoshimov et al. (2024); Chakhchi et al. (2024); Liu et al. (2024a, b); Khodadi et al. (2024); Liu et al. (2024c); Nojiri and Odintsov (2024). As a result, the possibility of PBHs accounting for the entire DM budget is severely constrained by a wide range of considerations and constraints. The only (not entirely undebated) remaining open window of parameter space where PBHs could make up all the DM is the so-called “asteroid mass window”, roughly for PBH masses $10^{14}\,{\text{kg}}\lesssim M_{\text{pbh}}\lesssim 10^{20}\,{\text{kg}}$ Katz et al. (2018); Smyth et al. (2020); Coogan et al. (2021); Ray et al. (2021); Auffinger (2022); Ghosh and Mishra (2024); Miller et al. (2022); Branco et al. (2023); Bertrand et al. (2023); Tran et al. (2023); Gorton and Green (2024); Dent et al. (2024); Tamta et al. (2024); Tinyakov (2024); Loeb (2024): lighter PBHs would have evaporated fast enough to either have disappeared by now or overproduced $\gamma$-rays in the MeV range, whereas heavier PBHs would have been detected through the microlensing of background stars.
@@
Almost all works on PBHs assume that these are Schwarzschild or Kerr BHs Khlopov (2010); Carr et al. (2016a); Green and Kavanagh (2021); Carr and Kuhnel (2020); Villanueva-Domingo et al. (2021); Carr and Kuhnel (2022); Bird et al. (2023); Carr et al. (2024); Arbey (2024); Choudhury and Sami (2024). All constraints and considerations on DM potentially being in the form of PBHs are therefore subject to this underlying assumption. The existence of the asteroid mass window, and the extension thereof, is of course no exception. The assumption in question is not at all unreasonable at least from the phenomenological point of view, given that there are at present no signs of tension between astrophysical observations and the Kerr-Newman family of metrics, and more generally the no-hair theorem. Nevertheless, from the theoretical point of view such an assumption might stir some unease, given the appearance of singularities in the Schwarzschild and Kerr metrics. The above considerations naturally lead to the following question: “what if PBHs are non-singular”? The present work is a pilot study whose goal is to explore this question, which naturally merges the DM and singularity problems.
@@
We entertain the possibility that PBHs are “regular”, i.e. free of curvature singularities Ansoldi (2008); Nicolini (2009); Sebastiani and Zerbini (2022); Torres (2022); Lan et al. (2023), and therefore that DM may be in the form of primordial regular BHs (PRBHs). For concreteness, we consider so-called tr (time-radius)-symmetric metrics, for which the product of the coefficients for the $dt^{2}$ and $dr^{2}$ components of the line element in four-dimensional Boyer–Lindquist coordinates is equal to $-1$, and the function which multiplies the angular part of the line element is $r^{2}$, i.e. $r$ is the areal radius. More specifically, we focus on the following four regular, static spherically symmetric space-times, all of which are characterized by an additional regularizing parameter $\ell$ and recover the Schwarzschild space-time in the $\ell\to 0$ limit: Bardeen BHs Bardeen (1968), Hayward BHs Hayward (2006), Culetu-Ghosh-Simpson-Visser BHs Culetu (2013, 2015); Ghosh (2015); Simpson and Visser (2019), and a Hayward-like BH which, to the best of our knowledge, is introduced for the first time in this work. These four space-times present a rich phenomenology, including either de Sitter or Minkowski cores, and temperatures which can both decrease and increase with increasing regularizing parameter $\ell$. We focus our attention on observational constraints from PRBH evaporation, which set the lower edge of the asteroid mass window, discussing in detail how the evaporation process is modified with respect to that for Schwarzschild PBHs. We show that, as a result, the phenomenology of PRBHs can be quite different from that of Schwarzschild PBHs, with a larger range of masses where PRBHs could make up the entire DM component, opening up the asteroid mass window by up to an extra decade in mass. Keeping in mind that the metrics in question are phenomenological in nature, our results demonstrate that a common solution to the DM and singularity problems in the form of PRBHs is one which is worth taking seriously, and warrants further investigation, and more generally the interface of these two problems provides a promising arena. We stress that our work should not be intended as a comprehensive analysis of PRBHs, but rather as a pilot study, pointing towards a direction which has thus far received little attention and indicating promising directions for further work.
@@
The rest of this paper is then organized as follows. In Sec. II we briefly introduce the regular space-times studied in the rest of the work. Various aspects of our methodology are discussed in Sec. III, with Sec. III.1 devoted to the calculation of the so-called greybody factors, Sec. III.2 to the computation of photon spectra resulting from Hawking evaporation, and Sec. III.3 to the comparison against observations. The resulting limits on the fraction of DM which may be in the form of primordial regular BHs are then critically discussed in Sec. IV. Finally, in Sec. V we draw concluding remarks. A number of more technical aspects concerning the greybody factors computation are discussed in Appendix A. Unless otherwise specified, we adopt units where $G=c=\hbar=1$. In closing, we note that a related study is being presented in a companion paper Calzà et al. (2024a): this focuses on non-tr-symmetric metrics, which also include loop quantum gravity-inspired metrics, but at the same time complicate the study of the evaporation process. We recommend that the interested reader go through the present work prior to consulting our companion paper Calzà et al. (2024a).
@@ II Regular black holes
It is well known, thanks to the Penrose-Hawking singularity theorems, that continuous gravitational collapse in GR sourced by matter contents satisfying reasonable energy conditions leads to the appearance of singularities Penrose (1965); Hawking and Penrose (1970). These are regions of space-time where curvature invariants, i.e. sets of independent scalars constructed from the Riemann tensor and the metric, diverge (with the archetypal example being the central singularity in the Kerr-Newman family of metrics). These singularities are arguably unsatisfactory as they lead to a potential breakdown in predictivity (see Refs. Sachs et al. (2021); Ashtekar et al. (2021, 2022) for a different viewpoint). For this reason, they are oftentimes regarded as a manifestation of our lack of knowledge of (new) physics in the high-energy/high-curvature regime. A widespread belief is that quantum gravity effects on these scales would ultimately cure the singularity problem (and potentially lead to observable effects), although this is more of a hope supported only by a few first-principles studies Dymnikova (1992); Dymnikova and Galaktionov (2005); Ashtekar and Bojowald (2005); Bebronne and Tinyakov (2009); Modesto et al. (2011); Spallucci and Ansoldi (2011); Perez (2015); Colléaux et al. (2018); Nicolini et al. (2019); Bosma et al. (2019); Jusufi (2023a); Olmo and Rubiera-Garcia (2022); Jusufi (2023b); Ashtekar et al. (2023); Nicolini (2023).
@@
Even in the absence of a widely agreed upon theory of quantum gravity, one can still hope to make progress in understanding and taming singularities, while also potentially gaining intuition about the possible features of such a theory, through a more phenomenological approach. Under the assumption that a metric description maintains its validity, one can introduce metrics which are free of singularities in the entire space-time, and describe so-called regular BHs (RBHs) Bambi (2023). It is often (albeit not necessarily always) the case that RBH metrics are controlled by an extra parameter, which in what follows we shall refer to as regularizing parameter (and denote by $\ell$), typically recovering the Schwarzschild metric (for non-rotating RBHs) in the limit $\ell\to 0$. Several RBH metrics have been studied over the past decades, see e.g. Refs. Borde (1997); Ayon-Beato and Garcia (1998, 1999); Bronnikov and Fabris (2006); Berej et al. (2006); Bronnikov et al. (2012); Rinaldi (2012); Stuchlík and Schee (2014); Schee and Stuchlik (2015); Johannsen (2013); Myrzakulov et al. (2016); Fan and Wang (2016); Sebastiani et al. (2017); Toshmatov et al. (2017); Chinaglia and Zerbini (2017); Frolov (2018); Bertipagani et al. (2021); Nashed and Saridakis (2022); Simpson and Visser (2022); Franzin et al. (2022); Chataignier et al. (2023); Ghosh et al. (2023); Khodadi and Pourkhodabakhshi (2022); Farrah et al. (2023); Fontana and Rinaldi (2023); Boshkayev et al. (2023); Luongo et al. (2023); Luongo and Quevedo (2024); Cadoni et al. (2023); Giambò and Luongo (2024); Cadoni et al. (2024); Luongo et al. (2024); Sajadi et al. (2024); Javed et al. (2024); Ditta et al. (2024); Al-Badawi et al. (2024); Övgün et al. (2024); Corona et al. (2024); Bueno et al. (2024); Konoplya and Zhidenko (2024a); Pedrotti and Vagnozzi (2024); Bronnikov (2024); Kurmanov et al. (2024); Bolokhov et al. (2024); Agrawal et al. (2024); Belfiglio et al. (2024); Stashko (2024); Faraoni and Rinaldi (2024); Konoplya and Stashko (2024); Khodadi and Firouzjaee (2024); Calzà et al. (2024b) for an inevitably incomplete selection of examples, as well as Refs. Ansoldi (2008); Nicolini (2009); Sebastiani and Zerbini (2022); Torres (2022); Lan et al. (2023) for recent reviews on the subject.  [^1] Another interesting possibility are gravastars, which are not RBHs in a strict sense Mazur and Mottola (2023, 2004); Mottola (2023). While most of these metrics have been introduced on purely phenomenological grounds it is known that possible sources for several RBH metrics lie in theories of non-linear electrodynamics Bronnikov (2018); Ghosh and Walia (2021); Bronnikov and Walia (2022); Bronnikov (2022); Bokulić et al. (2023).
@@
As alluded to earlier, our interest in this work is to explore the possibility that primordial RBHs may play the role of DM. As a proof-of-principle in this sense we will establish constraints on $f_{\text{pbh}}$, the fraction of DM in the form of PRBHs, focusing on the asteroid mass window, whose extent we will show can be either opened or tightened depending on the features of the RBH metric. To the best of our knowledge, we are aware of only five works in this little explored direction Dymnikova and Khlopov (2015); Pacheco (2018); Arbey and Auffinger (2021); Arbey et al. (2022); Banerjee et al. (2024). Ref. Dymnikova and Khlopov (2015) studied primordial BHs with de Sitter interiors as DM candidates, but considering the case where the DM is actually constituted of remnants from the evaporation process. Ref. Pacheco (2018) studied the thermodynamics of primordial regular BHs, focusing however on the case where they do not evaporate, and therefore did not study constraints on $f_{\text{pbh}}$. Ref. Arbey and Auffinger (2021) studied the evaporation of a loop quantum gravity-inspired BH, and weakened constraints on $f_{\text{pbh}}$ were reported in a later proceeding Arbey et al. (2022) (which however does not appear to be widely known). Finally, Ref. Banerjee et al. (2024) studied signatures of primordial BHs with magnetic charge, which could be (as is often but necessarily the case) regular. The aim of this pilot study and our companion paper Calzà et al. (2024a) is instead to provide a more comprehensive investigation of primordial regular BHs, considering a more diverse set of metrics and investigating constraints on $f_{\text{pbh}}$ in detail.
@@
In addition, as stated earlier, we require our space-times to be tr-symmetric (the non-tr-symmetric case is covered in a companion paper Calzà et al. (2024a)), which imposes the following additional conditions:
@@
$ds^{2}=-f(r)dt^{2}+f(r)^{-1}dr^{2}+r^{2}\left[d\theta^{2}+\sin^{2}(\theta)d\phi^{2}\right]\,.$ (4)
@@
with the first equality following from the choice of focusing on tr-symmetric space-times. In the case of Schwarzschild BHs, where the metric function is $f(r)=1-2M/r$, one recovers the well-known expressions $r_{H}=2M$ and $T_{\text{Sch}}=1/8\pi M$. However, in more general space-times the horizon radius in Eq. (7) is not guaranteed to have a closed form expression, and the same therefore holds for the temperature in Eq. (5). In Fig. 1 we show the evolution of the temperatures (normalized to the temperature of Schwarzschild BHs, $T_{\text{Sch}}=1/8\pi M$) of the four regular space-times we will introduce shortly, as a function of the regularizing parameter $\ell$ (itself normalized to the event horizon radius $r_{H}$). As the Figure clearly shows, for three of these space-times (Bardeen, Hayward, and Culetu-Ghosh-Simpson-Visser) the temperature is a monotonically decreasing function of $\ell$, whereas for the fourth (Hayward-like), after a small initial dip, the temperature grows monotonically with growing $\ell$. As we will see, these different behaviours turn out to have important phenomenological consequences.